please help me simplify 5x + 7y - 3x + 2y
Answer by Theo(2978)

Your expression is:
5x + 7y - 3x + 2y
You simplify this by combining like terms.
Combining like terms means to perform the indicated operation on them.
5x and -3x are like terms.
The indicated operation is addition
5x + (-3x) = 5x - 3x = 2x
-----
7y and 2y are like terms.
The indicated operation is addition.
7y + 2y = 9y
Your simplified expression is 2x + 9y after you combined like terms.
-----
To confirm that you did this correctly, pick any values for x and y and replace them in the original equation and the combined equation to see if you get the same answer.
-----
take x = 2, y = 3
-----
Your original equation of 5x + 7y - 3x + 2y becomes 5*2 + 7*3 - 3*2 + 2*3
This becomes 10 + 21 - 6 + 6 which becomes 31
-----
Your combined equation of 2x + 9y becomes 2*2 + 9*3 = 4 + 27 = 31
-----
You get the same answer so you simplified the original expression accurately.
-----
Your original equation is 5x + 7y - 3x + 2y
Your simplified equation is 2x + 9y
-----